OpenStudy (cutiecomittee123):

The value of pi/4 is a solution to the equation 3 sqrt 2 sec theta+7 = 1

OpenStudy (cutiecomittee123):

this is true or false. thus far I tried to solve the equation and I got sec theta = sqrt 2/2

zepdrix (zepdrix):

sec theta = sqrt2/2 ? Hmm that doesn't look quite right :o

zepdrix (zepdrix):

\[\large\rm 3\sqrt2 ~\sec \theta+7=1\]Subtracting 7 gives,\[\large\rm 3\sqrt2 ~\sec \theta=-6\]Dividing by 3sqrt2 gives,\[\large\rm \sec \theta=\frac{-6}{3\sqrt2}\qquad\to\qquad \sec \theta=-\frac{2}{\sqrt2}\]Something more like that, ya? :)

OpenStudy (cutiecomittee123):

I took away 1 on each side. Then I divided both sides by 3 then by root 2 so actually in this case it would be sec theta = 2/sqrt2

OpenStudy (cutiecomittee123):

Oh yeah I see where that is the better answer and the negative shouldve carried over in my answer. okay

zepdrix (zepdrix):

Divide 2 by sqrt2 to get,\[\large\rm \sec \theta=-\sqrt 2\] Wait wait wait, why all this fancy business? Why not just plug pi/4 into the equation and see if it holds true? :)

OpenStudy (cutiecomittee123):

very true. let me try that. thanks for pointing that out!

OpenStudy (cutiecomittee123):

wait pi/4 replaces theta right?

zepdrix (zepdrix):

yes

OpenStudy (cutiecomittee123):

also i dont know to to enter sec into my calculator, there is no sec button

zepdrix (zepdrix):

\[\large\rm 3\sqrt2 ~\color{orangered}{\sec \frac{\pi}{4}}+7=1\] \[\large\rm 3\sqrt2 ~\color{orangered}{\frac{1}{\cos\frac{\pi}{4}}}+7=1\]You don't need to use a calculator anyway you silly billy! >.< Don't you remember your unit circle and those special values? :)

OpenStudy (cutiecomittee123):

Yeah I forgot that sec is 1/cos

OpenStudy (cutiecomittee123):

but I dont see how you can solve without a calculator

zepdrix (zepdrix):

You haven't learned special values yet? :o cosine of pi/4 = sqrt(2)/2, yes?

OpenStudy (cutiecomittee123):

yes i know this. itis on the unit circle. I am not able to do those things without using the unit circle yet

OpenStudy (cutiecomittee123):

so where ever cos is equal to pi/4 which is sqrt2/2

OpenStudy (cutiecomittee123):

therefore it is a solution

zepdrix (zepdrix):

Woah woah, hold on. You're jumping too far ahead. \[\large\rm 3\sqrt2 ~\color{orangered}{\frac{1}{\cos\frac{\pi}{4}}}+7=1\]Yes, cos(pi/4) is sqrt2/2, \[\large\rm 3\sqrt2 ~\color{orangered}{\frac{1}{\left(\frac{\sqrt2}{2}\right)}}+7=1\]

zepdrix (zepdrix):

Now simplify further. Remember how to deal with `division by a fraction`?

OpenStudy (cutiecomittee123):

No fractions are not my forte at all

zepdrix (zepdrix):

Remember your uhh, what is it that they teach you rascals nowadays.. um Keep Change Flip? :)

OpenStudy (cutiecomittee123):

I think I know that saying, yeah

zepdrix (zepdrix):

\[\large\rm \frac{1}{\sqrt2/2}=1\cdot\frac{2}{\sqrt2}\] `Keep` the numerator the same `Change` the operation from division to multiplication `Flip` the bottom fraction.

zepdrix (zepdrix):

So our orange mess turns into this:\[\large\rm 3\sqrt2 \cdot\color{orangered}{\frac{2}{\sqrt2}}+7=1\]

OpenStudy (cutiecomittee123):

I see, that wasnt so bad but it will probably take time for me to catch on. now what though?

zepdrix (zepdrix):

Keep simplifying, see whether or not the equality holds true.

zepdrix (zepdrix):

I would probably subtract the 7 again,\[\large\rm 3\sqrt2\cdot\frac{2}{\sqrt2}=-6\]And then simplify the left side to see if this holds true.

OpenStudy (cutiecomittee123):

okay i get 2/sqrt2=2/sqrt2

OpenStudy (cutiecomittee123):

i mean 2/sqrt2=-2/sqrt2

zepdrix (zepdrix):

Which is NOT true, ya? :)

OpenStudy (cutiecomittee123):

yeah so this means the answer is false that pi/4 is not a solution

zepdrix (zepdrix):

Yay good job \c:/ It turns out that 3pi/4 and 5pi/4 are solutions, but not pi/4.

OpenStudy (cutiecomittee123):

Cool thanks for your help. I will probably have a few more questions tonight so if you'd like to join dont hesitate. I am taking a test and i get stuck sometimes

zepdrix (zepdrix):

cool ୧ʕ•̀ᴥ•́ʔ୨